# A Review on Single-Problem Multi-Attempt Heuristic Optimization

### v1

@@ 4.1. Algorithm Selection
Algorithm Selection (AS) (Kotthoff, 2016) is motivated by the efficient allocation of computational resources between a set of heuristic algorithms that perform differently on different optimization problems. This set is also identified as Algorithm Portfolio (Souravlias et al., 2021), where it is argued that the resources are divided among its constituents. Consequently, AS iteratively specifies the heuristic algorithm in which to spend more resources.
An instantiation of our abstraction in AS topic is provided by Souravlias et al. (2019) to solve the combinatorial problem called the detection of circulant weighing matrices (Arasu and Gutman, 2010), using a portfolio consisting of two variants of the algorithms Tabu Search (Glover and Laguna, 1998), two variants of Variable Neighbourhood Search (Mladenović and Hansen, 1997), and three variants of Multistart Local Search (Lourenço et al., 2019), under different parameter configurations. In each iteration, the number of available CPUs is distributed among the portfolio constituents proportionally to their performances in previous executions, to execute again in parallel new initializations of each one for the same maximum number of evaluations. Consequently, in this case, the algorithm, parameter configuration and initialization are the components that differentiate the alternatives. Moreover, running different times multiple algorithms also involves the availability of multiple attempts. Finally, the new qualities observed when running new initializations of each algorithm and configuration pair allow updating the estimated performance for each pair and associating proportionally the number of new initializations to execute next.
@@ 4.2. Parameter Tuning
Parameter Tuning (PT) (Huang et al., 2020a) focuses on automating the parameter setting of heuristic algorithms, to facilitate non-expert users in identifying well-performing configurations and avoiding costly brute force techniques, such as grid search. This topic is closely related to evolutionary computation (Larrañaga, 2002; Glover et al., 2003), where Genetic Algorithms or Estimation of Distribution Algorithms, although originally proposed as efficient, easy to use and applicable to a wide range of problems (Holland, 2010), their effectiveness is extremely dependent on the correct selection of parameters such as population size, crossover probabilities and selection ratios (Harik and Lobo, 1999). This requires a tuning process, in which different configurations are tested to find the most appropriate one for the problem of interest. In machine learning, the topic is known as Hyperparameter Optimization (HPO) (Schede et al., 2022), where the objective is to find the hyperparameter configuration of the learning algorithm of a model, which after being learned on a training data set has the minimum loss function value on a test data set. Therefore, since it is a priori unknown which hyperparameter configuration will perform best on the training data set, the computational resources must be distributed among several learning executions carried out with different configurations. In short, PT determines in which parameter configuration to invest resources at each iteration.
@@ 4.3. Multi-Start
Multi-Start (MS) (Martí et al., 2019; Locatelli and Schoen, 2016) is a technique used to improve the performance of heuristic algorithms, especially in non-convex problems where multiple local optima may exist. It alternates the execution of a single heuristic algorithm with its restart at another location in the solution space. Each initialization defines an independent optimization process, potentially leading to different local optima. Therefore, MS defines the new location to be considered, which determines the unknown performance of the algorithm until it is executed again.
@@ 4.4. Resource Allocation
Resource Allocation (RA) focuses on speeding up the evaluation of a parameter configuration or initialization of a heuristic algorithm. This is done by gradually allocating more resources to the most promising configurations or initializations and discarding the poor ones early. When different hyperparameter configurations of the same learning algorithm are evaluated, this approach is also identified as Multi-fidelity for HPO (Bischl et al., 2023; Elshawi et al., 2019; Yang and Shami, 2020). A monotonically increasing computational cost is considered for the model fitting, identifying higher fidelity evaluations with longer execution times in fits. The fidelity can be measured in terms of training epoch, number of iterations or train set size used in the previous ones for the learning algorithm. The last case carries out a Progressive sampling (Provost et al., 1999), where the sample of the entire available train set is considered and its size is progressively increased to retrain learning algorithms that have shown better results in previous trainings with smaller samples. In contrast, when evaluating different initializations of the same algorithm, the approach is related to MS and the resources are measured by the number of visited solutions. Therefore, allocating resources dynamically allows identifying unpromising alternatives quickly to have as much time as possible for those that seem to be promising according to the observed results. Thus, RA involves determining the amount of resources to be considered in each iteration.
@@ 5. Taxonomy of SIMHO strategies
In the following, we define a taxonomy with three linked classification rules associated with the three previous items. Each section defines each rule in order. In them, we determine the categories considered, illustrate each category with examples, and finally classify all the reviewed works accordingly. Overall, this section provides a review that characterizes existing selection strategies abstracting from the component. In fact, we will show how the same selection strategies sometimes appear in different topics to select different components.
Figure 1. The first box (named Single-problem Multi-attempt Heuristic Optimization framework…) highlights in yellow the contributions of a review carried out throughout the proposed framework. The second box under it (named …together with a taxonomy of selection strategies) enumerates in purple the linked classification rules in blue of a SIMHO strategy, with the categories defined for each of them in green, and the contributions implied by the resulting taxonomy in yellow.
@@ 5.1. Updating criteria
Fixed. The component distribution is fixed when each component value is always assigned the same probability from iteration to iteration. Thus, the distribution remains static independently of how the observed information is extended (see the first column in Rule 1 of Figure 1). In the works where the set of parameters $H$ and initializations $\Xi$ define different alternatives (Cicirello and Smith, 2005; Souravlias et al., 2019; Parsopoulos et al., 2022; Doerr and Zheng, 2020a; Auger and Hansen, 2005; El-Mihoub et al., 2022), the distribution of initializations $p_{\Xi}$ is fixed. To extend the observed information of one of the available algorithms, it is executed again considering a randomly selected initialization, i.e., an initialization sampled from a static uniform distribution defined over $X$.
@@ 5.2. Quality meaning of probability in adaptive criterion
Table 3. Classification of the papers depending on the criterion used to update each component distribution and the quality meaning of probability derived from the adaptive ones.
@@
$p_{H}$ $p_{\Theta}$ $p_{\Xi}$ $p_{C}$ EMQ EQI
(Mathesen et al., 2021) (Luersen and Le Riche, 2004) (Voglis and Lagaris, 2009) (Meng et al., 2023) (Wang et al., 2023) - - Adaptive -
@@
(Li et al., 2018) - Adaptive - Scheduled
(Falkner et al., 2018) (Huang et al., 2020b) - Adaptive - Scheduled
(Huang et al., 2019) - Adaptive - Adaptive
(Klein et al., 2017) - Adaptive - Adaptive
(Schmidt et al., 2020) Adaptive Adaptive - -
(Thornton et al., 2013) Adaptive Adaptive - -
(Li et al., 2020) Adaptive Adaptive - -
(Das Dores et al., 2018) Adaptive Adaptive - Scheduled
(Hu et al., 2022) Adaptive Fixed Adaptive - -
@@ 5.3.1. Probability related to expected maximum quality
Table 5. Classification of papers depending on the meaning of the utility predictor and the observed information used for it.
Paper Utility predictor Which to use from $\mathcal{I}$?
$x_{a}$ $t_{a}$ $f_{a}$
(Souravlias et al., 2019) (Li et al., 2018; Das Dores et al., 2018) (György and Kocsis, 2011) (Parsopoulos et al., 2022) (Jaderberg et al., 2017) (Falkner et al., 2018) (Huang et al., 2020b) (Vrugt et al., 2009) (Siddiqui et al., 2025) Point estimation
(Hu et al., 2022)
(Huang et al., 2019) (Li et al., 2020) Interval estimation
(Mathesen et al., 2021) (Meng et al., 2023) (Cicirello and Smith, 2005) (Klein et al., 2017) (Thornton et al., 2013) (Addis et al., 2005) (Falkner et al., 2018) (Huang et al., 2020b) Density estimation
(Schmidt et al., 2020) (Hu et al., 2022) Extrapolation
@@
Table 6. Classification of the papers depending on how they control the uncertainty of utilities.
@@
(Schmidt et al., 2020) (Hu et al., 2022) (György and Kocsis, 2011) UCB + argmax
(Schmidt et al., 2020) $\varepsilon$-greedy
@@
(Mathesen et al., 2021) (Meng et al., 2023) (Klein et al., 2017) (Thornton et al., 2013) (Falkner et al., 2018) (Huang et al., 2020b) Acquisition function + argmax
(Luersen and Le Riche, 2004) (Jaderberg et al., 2017) (Voglis and Lagaris, 2009) (Addis et al., 2005) (Vrugt et al., 2009) (Wang et al., 2023) (Siddiqui et al., 2025) Random sampling
@@ 5.3.2. Probability related to expected quality improvement
The second option for guiding sampling toward the optimal alternative consists of associating higher probabilities to the component values expected to define alternatives with more chances for improving the best observed quality. For a component distribution that defines probabilities with this meaning, we found works that continue defining the utility predictor considering the qualities of the observed information (the $f_{a}\in\mathcal{I}$), although we also observed the use of only the best-visited solutions to predict utilities (the $x_{a}\in\mathcal{I}$). However, in all of them, it is always considered an implicit control of uncertainties.
@@
  • Utility prediction with $f_{a}$. The literature techniques that predict the utilities using the qualities correspond to the first stage of Bayesian optimization, i.e., the prediction by a model. Utility predictors are models that capture the dependence between the component value and the $f_{a}$ of the alternatives they define. The Gaussian process is the most commonly used surrogate model, which estimates a distribution for the quality of each unexecuted parameter configuration (Klein et al., 2017; Hu et al., 2022) or initialization (Mathesen et al., 2021; Meng et al., 2023) or stopping criterion (Klein et al., 2017), after training the model with the pairs formed by the executed component values and their observed qualities. Thornton et al. (2013) instead, use random forest models to predict simultaneously the quality of algorithm and parameter configuration combinations, arguing that they are good predictors with discrete and high-dimension input data. In other works (Falkner et al., 2018; Huang et al., 2020b), an adaptation of the Tree Parzen Estimator (Bergstra et al., 2011) is used to model densities over the executed parameter configuration space.
@@
  • Utility prediction with $x_{a}$. As an example of utility prediction that uses the best-visited solutions, Luersen and Le Riche (2004) estimate the density of the set formed by the pairs of starting points $\xi$ and local convergence points $x_{a}$ visited in each previous execution of the Nelder-Mead algorithm. They argue that the topology of the basins of attraction can be efficiently summarized by using only these two points among all the points visited during the local searches. The utility predictor is the density estimator obtained by applying a Gaussian Parzen-Window. Thus, the predicted utility for each new initialization is the probability that the Nelder-Mead departed from that point can explore a different region of the solution space and converge to a new local optimum. Similarly, Voglis and Lagaris (2009) estimate the probability that an initial point does not belong to the region of attraction of any of the previously identified minima. These probabilities depend on the distances between the initial points and the local minima found by the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm (Rinnooy Kan and Timmer, 1987), the direction of the gradient, and the frequency with which each minimum has been discovered. Wang et al. (2023) consider a similar potential metric for evolutionary computation methods. They show how the population of evolutionary methods with simple restarts tend to repeat previous evolution processes ending in the basins of attraction already explored. To address this issue, they use the evaluated solutions to divide the search space into hill regions for estimating the basins of attractions, and encourage further exploration in those of higher potential for finding unidentified ones.
@@
    When using a Bayesian optimization, the uncertainty of the quality distributions predicted by the model for each component value is controlled by an acquisition function. Among the most popular options is the expected improvement (Mathesen et al., 2021; Meng et al., 2023; Falkner et al., 2018; Huang et al., 2020b). It maps to each component value the expected quality increment over the best observed, which is larger for values predicted with higher mean or variance of quality. Klein et al. (2017) also use the more recent entropy search acquisition function. It favours component values with larger expected information gains about the optimum, rather than trying to evaluate close to the optimum. Notice that the larger the observed information, the predicted distributions have a lower variance, leading to closer acquisition functions to the predicted model. Thus, the number of iterations implicitly measures the uncertainty.
@@
6. Framework utility and design of new strategies

### v2

@@ 4.1. Algorithm Selection
Algorithm Selection (AS), also identified as Algorithm Portfolio by Souravlias et al. (2021), is motivated by the efficient allocation of computational resources between a set of heuristic algorithms that perform differently on the problem of interest. Consequently, AS iteratively specifies the heuristic algorithm in which to spend more resources.
An instantiation of our abstraction in the AS topic is provided by Souravlias et al. (2019) to solve the combinatorial problem called the detection of circulant weighing matrices (Arasu and Gutman, 2010), using a portfolio consisting of two variants of the algorithm Tabu Search (Glover and Laguna, 1998), two variants of Variable Neighbourhood Search (Mladenović and Hansen, 1997), and three variants of Multistart Local Search (Lourenço et al., 2019), under different parameter configurations. In each iteration, the number of available CPUs is distributed among the portfolio constituents proportionally to their performances in previous executions, to execute again in parallel new initializations of each one for the same maximum number of evaluations. Consequently, in this case, the algorithm, parameter configuration and initialization are the components that differentiate the alternatives. Moreover, running different times multiple algorithms also involves the availability of multiple attempts. Finally, the new qualities observed when running new initializations of each algorithm and configuration pair allow updating the estimated performance for each pair and proportionally associating the number of new initializations to execute next.
@@ 4.2. Parameter Tuning
Parameter Tuning (PT), denoted as specialist PT by Huang et al. (2020a), focuses on automating the parameter setting of heuristic algorithms, to facilitate non-expert users in identifying well-performing configurations to the problem of interest and avoiding costly brute force techniques, such as grid search. This topic is closely related to evolutionary computation (Larrañaga, 2002; Glover et al., 2003), where Genetic Algorithms or Estimation of Distribution Algorithms, although originally proposed as efficient, easy to use and applicable to a wide range of problems (Holland, 2010), their effectiveness is extremely dependent on the correct selection of parameters such as population size, crossover probabilities and selection ratios (Harik and Lobo, 1999). This requires a tuning process, in which different configurations are tested to find the most appropriate one for the problem of interest. In machine learning, the topic is known as Hyperparameter Optimization (HPO) (Bischl et al., 2023) where the objective is to find the hyperparameter configuration of the learning algorithm of a model, which after being learned on a training data set has the minimum loss function value on a test data set. We do not know a priori which hyperparameter configuration will perform best, and thus the computational resources must be distributed among several learning executions carried out with different configurations. In short, PT determines in which parameter configuration to invest resources at each iteration.
@@ 4.3. Multi-Start
Multi-Start (MS) (Martí et al., 2019; Locatelli and Schoen, 2016) is a technique used to improve the performance of heuristic algorithms, especially in non-convex problems where multiple local optima may exist. It alternates the execution of a single heuristic algorithm with its restart at another location in the solution space. Each initialization defines an independent optimization process, potentially leading to different local optima. In machine learning, HPO may require substantial computational resources, and one way to reduce the computational time are warm-starts, discussed by (Bischl et al., 2023) as MS strategies that use information from previous experiments to initialize new ones. Therefore, MS defines the new location to be considered, which determines the unknown performance of the algorithm until it is executed again.
@@ 4.4. Resource Allocation
Resource Allocation (RA) focuses on speeding up the evaluation of a parameter configuration or initialization of a heuristic algorithm. This is done by gradually allocating more resources to the most promising configurations or initializations and discarding the poor ones early. When different hyperparameter configurations of the same learning algorithm are evaluated, this approach is also identified as Multi-fidelity for HPO (Bischl et al., 2023; Elshawi et al., 2019; Yang and Shami, 2020). A monotonically increasing computational cost is considered for the model fitting, identifying higher fidelity evaluations with longer execution times in fits. The fidelity can be measured in terms of training epoch, number of iterations or train set size used in the previous ones for the learning algorithm. The last case carries out a Progressive sampling (Provost et al., 1999), where a sample of the entire available train set is considered and its size is progressively increased to retrain learning algorithms that have shown better results in previous trainings with smaller samples. In contrast, when evaluating different initializations of the same algorithm, the approach is related to MS and the resources are measured by the number of visited solutions. Therefore, dynamically allocating resources allows identifying unpromising alternatives quickly to have as much time as possible for those that seem to be promising according to the observed results. Thus, RA involves determining the amount of resources to be considered in each iteration.
@@ 5. Taxonomy of SIMHO strategies
Figure 1. The first box (named Single-problem Multi-attempt Heuristic Optimization framework…) highlights in yellow the contributions of a review carried out throughout the proposed framework. The second box under it (named …together with a taxonomy of selection strategies) enumerates in purple the linked classification rules in blue of a SIMHO strategy, with the categories defined for each of them in green, and the contribution implied by the resulting taxonomy in yellow.
@@
In the following, we define a taxonomy with three linked classification rules associated with the three previous items. Each section defines each rule in order. In them, we determine the categories considered, illustrate each category with examples, and finally classify all the reviewed works accordingly. Overall, this section provides a review that characterizes existing single-problem selection strategies abstracting from the component. In fact, we will show how the same selection strategies sometimes appear in different topics to select different components. The papers included in this section extend and complete the illustrative examples described in Section 4, all of them presenting strategies that we consider single-problem [^7] Some of the strategies reviewed in Section 5 are presented in the literature without explicitly defining the specific solution space and objective function of the optimization problem they aim to address. Nevertheless, these strategies remain inherently single-problem, as they rely exclusively on the information generated during the optimization of the same problem.. These strategies fit with the abstract Algorithm 1 that formalizes under the same terminology those valid for SIMHO. They constitute the core of our review [^8] The related reviews summarized in Table 1 are not classified under our taxonomy, as they are used only for comparison purposes and are not part of the reviewed works themselves., which are summarized and classified through Tables 3, 5, and 6.
@@ 5.1. Updating criteria
Fixed. The component distribution is fixed when each component value is always assigned the same probability from iteration to iteration. Thus, the distribution remains static independently of how the observed information is extended (see the first column in Rule 1 of Figure 1). In the works where the set of algorithms $H$ or parameters $\Theta$ and initializations $\Xi$ define different alternatives (Cicirello and Smith, 2005; Souravlias et al., 2019; Parsopoulos et al., 2022; Doerr and Zheng, 2020a; Auger and Hansen, 2005; El-Mihoub et al., 2022), the distribution of initializations $p_{\Xi}$ is fixed. To extend the observed information of one of the available algorithms or parameters, it is executed again considering a randomly selected initialization, i.e., an initialization sampled from a static uniform distribution defined over $X$.
@@ 5.2. Quality meaning of probability in adaptive criterion
Table 3. Classification of reviewed papers depending on the criterion used to update each component distribution and the quality meaning of probability derived from the adaptive ones. The classification of papers with at least one adaptive component is completed in the sections indicated in the last column and summarized in Tables 5 and 6.
@@
$p_{H}$ $p_{\Theta}$ $p_{\Xi}$ $p_{C}$ EMQ (Section 5.3.1) EQI (Section 5.3.2)
@@
(Luersen and Le Riche, 2004) (Mathesen et al., 2021) (Meng et al., 2023) (Voglis and Lagaris, 2009) (Wang et al., 2023) - - Adaptive -
(Thornton et al., 2013) Adaptive Adaptive - -
(Hu et al., 2022) Adaptive Fixed Adaptive - -
(Li et al., 2020) (Schmidt et al., 2020) Adaptive Adaptive - -
(Cicirello and Smith, 2005) Adaptive - Fixed -
(El-Mihoub et al., 2022) Fixed - Fixed -
@@
(Huang et al., 2019) - Adaptive - Adaptive
(Klein et al., 2017) - Adaptive - Adaptive
(Li et al., 2018) - Adaptive - Scheduled
(Falkner et al., 2018) (Huang et al., 2020b) - Adaptive - Scheduled
@@ 5.3.1. Probability related to expected maximum quality
Table 5. Classification of reviewed papers that address the selection of at least one adaptive component depending on the meaning of the utility predictor and the observed information used for it. Papers are grouped by common utility predictor and ordered following their description in Sections 5.3.1 and 5.3.2. A paper may appear multiple times when different utility predictors are considered for updating the adaptive components shown in Table 3.
Paper Utility predictor Which to use from $\mathcal{I}$?
$x_{a}$ $t_{a}$ $f_{a}$
(Das Dores et al., 2018) (Falkner et al., 2018) (György and Kocsis, 2011) (Huang et al., 2020b) (Jaderberg et al., 2017) (Li et al., 2018) (Parsopoulos et al., 2022) (Siddiqui et al., 2025) (Souravlias et al., 2019) (Vrugt et al., 2009) Point estimation
(Hu et al., 2022)
(Addis et al., 2005) (Cicirello and Smith, 2005) (Falkner et al., 2018) (Huang et al., 2020b) (Klein et al., 2017) (Mathesen et al., 2021) (Meng et al., 2023) Density estimation
(Huang et al., 2019) (Li et al., 2020) Interval estimation
(Thornton et al., 2013)
(Hu et al., 2022) (Schmidt et al., 2020) Extrapolation
@@
Table 6. Classification of reviewed papers that address the selection of at least one adaptive component depending on how they control the uncertainty of utilities. Papers are grouped by common uncertainty controller and ordered following their description in Sections 5.3.1 and 5.3.2. A paper may appear multiple times when different uncertainty controllers are considered for updating the adaptive components shown in Table 3.
@@
(György and Kocsis, 2011) (Hu et al., 2022) (Schmidt et al., 2020) UCB + argmax
(Das Dores et al., 2018) (Falkner et al., 2018) (Huang et al., 2020b) (Li et al., 2018) Successive discarding
@@
(Schmidt et al., 2020) $\varepsilon$-greedy
(Falkner et al., 2018) (Huang et al., 2020b) (Klein et al., 2017) (Mathesen et al., 2021) (Meng et al., 2023) (Thornton et al., 2013) Acquisition function + argmax
(Addis et al., 2005) (Jaderberg et al., 2017) (Luersen and Le Riche, 2004) (Siddiqui et al., 2025) (Voglis and Lagaris, 2009) (Vrugt et al., 2009) (Wang et al., 2023) Random sampling
(Cicirello and Smith, 2005) Boltzmann exploration
@@ 5.3.2. Probability related to expected quality improvement
The second option for guiding sampling toward the optimal alternative consists of associating higher probabilities to the component values expected to define alternatives with more chances for improving the best observed quality. For a component distribution that defines probabilities with this meaning, we found works that continue defining the utility predictor considering the qualities of the observed information (the $f_{a}\in\mathcal{I}$), although we also observed the use of only the best-visited solutions to predict utilities (the $x_{a}\in\mathcal{I}$). However, in all of them, an implicit control of uncertainties is always considered.
@@
  • Utility prediction with $f_{a}$. The literature techniques that predict the utilities using the qualities correspond to the first stage of Bayesian optimization, i.e., the prediction by a model. Utility predictors are models that capture the dependence between the component value and the $f_{a}$ of the alternatives they define. The Gaussian process is the most commonly used surrogate model, which estimates a distribution for the quality of each unexecuted parameter configuration (Klein et al., 2017; Hu et al., 2022) or initialization (Mathesen et al., 2021; Meng et al., 2023) or stopping criterion (Klein et al., 2017), after training the model with the pairs formed by the executed component values and their observed qualities. Thornton et al. (2013) instead, use random forest models to predict simultaneously the quality mean and variance of algorithm and parameter configuration combinations, arguing that they are good predictors with discrete and high-dimensional input data. In other works (Falkner et al., 2018; Huang et al., 2020b), an adaptation of the Tree Parzen Estimator (Bergstra et al., 2011) is used to model densities over the executed parameter configuration space.
@@
  • Utility prediction with $x_{a}$. As an example of utility prediction that uses the best-visited solutions, Luersen and Le Riche (2004) estimate the density of the set formed by the pairs of starting points $\xi$ and local convergence points $x_{a}$ visited in each previous execution of the Nelder-Mead algorithm. They argue that the topology of the basins of attraction can be efficiently summarized by using only these two points among all the points visited during the local searches. The utility predictor is the density estimator obtained by applying a Gaussian Parzen-Window. Thus, the predicted utility for each new initialization is the probability that the Nelder-Mead departed from that point can explore a different region of the solution space and converge to a new local optimum. Similarly, Voglis and Lagaris (2009) estimate the probability that an initial point does not belong to the region of attraction of any of the previously identified minima. These probabilities depend on the distances between the initial points and the local minima found by the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm (Rinnooy Kan and Timmer, 1987), the direction of the gradient, and the frequency with which each minimum has been discovered. Wang et al. (2023) consider a similar potential metric for evolutionary computation methods. They show how the populations of evolutionary methods with simple restarts tend to repeat previous evolution processes ending in the basins of attraction already explored. To address this issue, they use the evaluated solutions to divide the search space into hill regions for estimating the basins of attractions, and encourage further exploration in those of higher potential for finding unidentified ones.
@@
    When using a Bayesian optimization, the uncertainty of the quality distributions predicted by the model for each component value is controlled by an acquisition function. Among the most popular options is the expected improvement. It maps to each initialization (Mathesen et al., 2021; Meng et al., 2023) or parameter configuration (Falkner et al., 2018; Huang et al., 2020b) the expected quality increment over the best observed, which is larger for values predicted with higher mean or variance of quality. Klein et al. (2017) also use the more recent entropy search acquisition function. It favours component values with larger expected information gains about the optimum, rather than trying to evaluate close to the optimum. Notice that the larger the observed information, the predicted distributions have a lower variance, leading to acquisition functions closer to the predicted model. Thus, the number of iterations implicitly measures the uncertainty.
@@
6. Discussion
This section discusses the implications of the proposed framework and the scope of the review. First, we reflect on the utility of the framework as a conceptual tool to organize existing single-problem selection strategies and to support the design of new ones. Second, we discuss the relationship between single- and multi-problem settings, with a particular focus on the transferability of strategies across these settings and on the intended scope and limitations of this review.
6.1. Framework utility and design of new strategies